A firm faces the following linear inverse demand for its product P = 60 - 2Q.
a) Find the firm's total revenue function TR (Q). (4 marks)
b) Find the expression for the firm's marginal revenue. (2 marks)
c) Assuming that the marginal cost of production is given by MC=8. What will be the equilibrium output and price? (4 marks)
1
Expert's answer
2021-07-23T05:22:27-0400

a) The firm's total revenue function is:"TR(Q) = P\u00d7Q = 60Q - 2Q^2."

b) The expression for the firm's marginal revenue is:

MR(Q) = TR'(Q) = 60 - 4Q.

c) Assuming that the marginal cost of production is given by MC = 8, the equilibrium output and price are:

60 - 4Q = 8,

4Q = 52,

Q = 13 units,

P = 60 - 2×13 = 34.